Can you please help me to solve with steps this type of problems?


Q:the 20th term of the series 2*4+4*6+6*8+........is


a)1600   b)1680   c)420   d)840.


 

Grade:11

2 Answers

manni stubtza
13 years ago

1st takes it general term means Nth term which is as seen from sequence is 2n[2n+2]

u may verify it.....

on put n=1,2.... we get our 1st and 2nd term and so on....

hence for 20th term put n=20

so 2*20[2+2*20]=1680 which is our ans.

hence correct option is B

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vivek

carefully look at the terms in each multiplication the two terms form  AP. like first terms are 2,4,6,8,.... with common difference 2 and first term being 2

20 th term= a+(n-1)d

               = 2+(20-1)2

              = 40

for the second term 4,6,8,..... common difference is 2 and first term being 4

using above formula

20th term= 4+(20-1)2 = 42

20 th term of given series = 40*42 = 1680

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..